Given that the point P(2, 6) lies on the line 4x + ky - 19 = 0, find k.
Everytime I do this problem, I get the answer k= 11/6, but this is not one of my choices. Am I correct in my answer?

Answer by Alan3354(69443)

Given that the point P(2, 6) lies on the line 4x + ky - 19 = 0, find k.
4x + ky = 19
4*2 + k*6 = 19
6k = 11
k = 11/6
-----------
Your answer is right.
------------
Everytime I do this problem, I get the answer k= 11/6, but this is not one of my choices. Am I correct in my answer?